# UniMS-RAG: A Unified Multi-source Retrieval-Augmented Generation for Personalized Dialogue Systems

### v1

@@ 5.3.3. Response Generation
(11) $\textbf{C}(r,g)=\begin{cases}\textbf{NLI}(r,g),&\text{if there is grounding $g$ for $r$, and planning also uses $g^{\prime}$.}\\ 0,&\text{if there is grounding $g$ for $r$, and planning {does not} use $g^{\prime}$.}\\ 0,&\text{if there is {no} grounding $g$ for $r$, and planning uses $g^{\prime}$.}\\ 1,&\text{if there is {no} grounding $g$ for $r$, and planning {does not} use $g^{\prime}$.}\\ \end{cases}$
5.4. Implementation Details
UniMS-RAG: We mainly choose ChatGLM-6B (Du et al., 2022; Zeng et al., 2023) as the backbone models during training, we set the batch size as 8, train models with 3 epochs and save the checkpoint with the lowest validation loss. For other hyper-parameter settings, we mainly follow the corresponding official code [^8] https://github.com/THUDM/ChatGLM-6B. Due to the computation limit, we conduct training with LoRA (Hu et al., 2021) at one single 3090 GPU, and it costs about 4-6 hours. For the prompting using ChatGPT, we choose gpt-3.5-turbo-1106 and set both the temperature and top p as 0.1 to reduce the randomness of LLMs. We set the number of retrieved results as 3 to evaluate whether or not the UniMS-RAG can denoise the unrelated evidence. According to different sources of similarity scores, we have two variants: UniMS-RAG (ChatGPT) and UniMS-RAG (DPR) where the similarity scores come from ChatGPT and DPR, respectively. At the inference stage, we use self-predicted similarity scores to retrieve the evidence, serving as an indicator. More variants or analyses can be found in §7.2.
Others: We finetune RocketQAv2 and DPR using samples from corresponding dataset by regarding (context, used_persona / document) as the positive and (context, unrelated_persona / document) as the negative. We set epochs as 5 and max sequence length as 512, and mainly follow these codebases [^9] https://github.com/PaddlePaddle/RocketQA, https://github.com/Alibaba-NLP/Multi-CPR/tree/main/retrieval for other parameters. For RocketQAv2, we load the weights of pre-trained model zh_dureader_de_v2 as introduced in the official homepage, which is trained on the largest Chinese QA dataset, and we use 12-layer bert-base-chinese with 110M parameters as backbone model for DPR. We then finetune an NLI model (Madotto et al., 2019) by regarding (ground_persona / document, response) as the positive and randomly sampled (unrelated_persona / document, response) as the negative. We also use bert-base-chinese as the backbone model. We concatenate and encode the ground persona/document $k$ and response $r$ in the form of $\mathrm{[CLS]}k\mathrm{[SEP]}r\mathrm{[SEP]}$, and we train the model to predict whether responses are consistent with corresponding personas or documents. The batch size for fine-tuning is 8. The maximal training epoch is 5, and the max sequence length of the encoder is 512. In the experiments, we use the AdamW optimizer with a learning rate of 2e-5 and an epsilon of 1e-6. We evaluate the NLI model on the KBP test set every 500 iterations during training, and we save the checkpoint with the highest performance on the test set. The fine-tuned NLI model achieves $>95\%$ accuracy for both datasets.
6. Experimental Results
6.1. Performance of Planning (RQ1)
Table 3. The F1 of different decisions in Planning of different methods under supervised settings. 1) DuLeMon: There are 1686 NULL, 505 USER, 219 BOT and 6 BOTH in the ground planning; 2) KBP: There are 181 NULL, 125 PERSONA and 923 BOTH in the ground planning.
@@ 6.1. Performance of Planning (RQ1)
There are different types of planning decisions in the different datasets: DuLeMon (using NULL, User-Per, Bot-Per and Both sources of knowledge) and KBP (using NULL, Persona, and Both). Table 3 demonstrates the F1 of planning using different methods under these two datasets. Specifically, we found that SAFARI is comparable with the BERT model, and furthermore UniMS-RAG achieves better performance than SAFARI on DuLeMon but slightly worse on KBP. In addition, we also found that the original data distribution has a serious impact on the final planning performance. For example, there are less than 0.1% samples in the training set of DuLeMon requiring Both source of knowledge, resulting in poor performance of all methods at Both. A similar phenomenon is also observed on the KBP dataset. Another reason behind this could be the additional token (i.e., evaluation tokens) introduced during the training stage compared with SAFARI. In general, the planning capability of LLMs still needs to be improved to solve the complex multiple sources planning problem in a dialogue system, particularly when dealing with imbalanced or scarce data sources.
Table 4. The performance (Recall@1) of Retrieval of different types of retrievers. It is note the UniMS-RAG (w/ DPR) means that the similarity labels come from DPR model during training, and UniMS-RAG (w/ LLM) means that the similarity labes come from ChatGPT model. 1) DuLeMon: There are 511 samples require to use User-Per and 225 samples require to use Bot-Per; 2) KBP: There are 125 samples require to use PERSONA and 923 samples require to use BOTH.
Model DuLeMon KBP
User-Per Bot-Per PERSONA Both-PERSONA Both-DOCUMENTS
BM25 15.85 24.00 36.80 48.97 15.05
RocketQAv2 42.85 46.22 80.00 92.31 50.49
DPR 30.72 40.00 83.20 93.07 51.67
ChatGPT 21.14 37.33 45.60 65.33 16.14
[…]
 w/ DPR 23.87 31.11 74.40 82.88 22.32
 w/ ChatGPT 21.53 33.33 42.40 67.61 14.41
6.2. Performance of Retrieval (RQ2)
To investigate the RQ2, we examine different types of retrievers, including BM25, RocketQAv2, DPR, ChatGPT and our proposed UniMS-RAG to evaluate the retrieval performance, providing the ground-truth planning labels (except NULL). Table 4 presents the Recall@1 (R@1) of the different retrievers.
6.2.1. The performance of baselines
There are several observations can be concluded from the results: 1) The overall performance of dense vector (e.g, RocketQAv2 and DPR) is better than ChatGPT, and the performance of ChatGPT is better than sparse retriever (e.g, BM25). The gap of former is bigger than the latter, indicating the large improvement room for current methods using LLMs as retriever, particularly at the context of conversational search; 2) It is observed that DPR performs the best out of these retrievers on KBP datasets while RocketQAv2 performs the best on DuLeMon dataset, revealing the importance of dense retrieval models in this task. We attribute the higher performance of RocketQAv2 on DuLeMon to the similar distribution between DuLeMon dataset and pre-training corpus of RocketQAv2; 3) On KBP datasets, all retrievers performs best at Both-PERSONA and worst at Both-DOCUMENTS, indicating the difficulty to retrieve independent source of knowledge since the semantics between different knowledge from Both-DOCUMENTS are similar to the same underlying persona $p^{*}$, making them more difficult to distinguish. On the other hand, the performance on DuLeMon is even worse than Both-DOCUMENTS on KBP. We suspect this is due to the semantic gap between the dialogue context and used persona in DuLeMon (as shown in Figure 1).
6.2.2. The performance of UniMS-RAG
Besides, we also present the performance of UniMS-RAG as a retriever using different similarity signals (DPR or ChatGPT) during the training. The results show that UniMS-RAG can be directly used as a retriever, since the performance of both UniMS-RAG w/ DPR and UniMS-RAG w/ ChatGPT are better than sparse retriever (BM25). In detail, UniMS-RAG w/ DPR even achieves better performance than ChatGPT, revealing the great potential of UniMS-RAG as retriever once we use more high-quality signals (the performance of DPR is better than ChatGPT). Furthermore, UniMS-RAG w/ ChatGPT achieves comparable performance with ChatGPT on DuLeMon and KBP datasets. Since the range of similarity signals provided by ChatGPT is much smaller than DPR, we believe that UniMS-RAG can distill the capabilities of original retriever as much as possible if we can provide more data and more fine-grained signals. Therefore, we can derive the answer to RQ2 from this analysis: Large Language Models (LLMs) can serve as retrievers directly, achieving comparable performance compared with original retrievers, showcasing a great potential towards a more powerful unified RAG framework.
Table 5. The performance of Generation of different methods, in which part of the results are directly copied from Wang et al. (2023a). We follow the official code of SAFARI to conduct evaluation on DuLeMon dataset.
Models DuLeMon KBP
BLEU-1 Rouge-L P.C BLEU-1 Rouge-L P.C K.C
Unsupervised Setting
SAFARI (Wang et al., 2023a) 12.11 15.46 4.96 13.74 19.69 16.92 24.89
Supervised Setting
FoCus (Jang et al., 2022b) - - - 22.51 25.29 64.52 17.90
SAFARI (Wang et al., 2023a) 7.81 8.72 51.18 23.81 26.70 76.99 42.39
Using itself as retriever
[…]
 w/ DPR 18.43 20.32 63.18 29.65 32.48 75.51 42.07
 w/ ChatGPT 18.61 20.33 64.83 27.86 30.86 70.38 36.53
Using independent retriever
[…]
 w/ BM25 18.52 20.34 64.07 27.30 29.22 68.99 36.29
 w/ DPR 18.95 20.87 64.28 29.72 32.85 73.56 45.00
 w/ ChatGPT 18.53 20.47 65.56 29.04 31.46 72.98 39.62
6.3. Performance of Generation (RQ3)
@@
To investigate the performance of UniMS-RAG on the response generation task, we conduct two settings: 1) using itself as retriever, which means we use UniMS-RAG to retrieve corresponding evidences from the planned sources of knowledge (i.e, the output of planning step); and 2) using independent retriever (i.e, BM25, DPR, ChatGPT) to retrieve corresponding evidences from the planned sources of knowledge (i.e, the output of planning step). Table 5 demonstrates the performance of response generation under both supervised and unsupervised settings.
6.3.1. The performance of baselines
@@ 6.3.2. The performance of UniMS-RAG
Referring to Tabel 3 and Table 4, the performance of the planning step and retrieval step largely affects the results in the final generation step. Specifically, when using itself as retriever, we can find that UniMS-RAG using signals from DPR (w/ DPR) leads to better performance in contrast to ChatGPT (w/ ChatGPT), revealing the effectiveness of better retriever signals. The gap between DPR and ChatGPT on DuLeMon is relatively small since most of cases here do not require the involvement of external sources of knowledge. Thus, we decide to load parameters of UniMS-RAG w/ DPR to conduct evaluation when using independent retriever. In detail, the results again validate the effectiveness of better retriever (w/ DPR $>$ w/ ChatGPT $>$ w/ BM25). We find that the performance gap between different retrievers in DuLeMon dataset is pretty small (less than 1%). We suspect this is highly related to the original data distribution in the test dataset, since most of the samples do not require external persona information. Furthermore, we also find that using independent retriever is mostly better than using itself as retriever except the worst BM25, which is consistent with the findings in the performance of retrieval step. To conclude, it is obvious that our proposed method UniMS-RAG outperforms all other baselines in at least 5 out 7 evaluation metrics no matter using itself as retriever or using independent retriever. Combining the performance of planning, retrieval, and response generation together, we can find that UniMS-RAG is capable of serving as a planner, retriever, and reader in a unified manner, leading to better performance in personalized dialogues.
7. Analysis and Discussions
@@
In this section, we choose the best model as shown in our previous experiments to investigate the performance changes under different settings (UniMS-RAG w/ DPR). In detail, we start from the performance of our proposed model with different numbers of retrieved evidence (§7.1), then we study the effects of introduced self-refinement mechanism during the inference stage by re-evaluating the relationship between generated response with the dialogue context or retrieved evidence (§7.2). We present the ablation study to show the effectiveness and rationale of UniMS-RAG (§7.2), followed by the results of human evaluation (§7.2). Finally, We conclude this section with case study (§7.5) and error analysis (§7.6).
7.1. Different Numbers of Retrieved Results
@@
(a) Performance on DuLeMon dataset.
(b) Performance on KBP dataset.
Figure 4. The performance of Generation with different number of retrieved evidences on two datasets.
7.2. Additional Self-refinement during Inference
To investigate the effects of self-refinement during the inferences, we set the number of retrieved evidence as 3 according to the experimental results in the previous section since it leads to the best performance. There are two different ways to refine the generated response: 1) Within self-refinement by providing different number of updated evidences during one-step refinement as shown in Algorithm 1; and 2) Multi-step self-refinement by refining the response step by step while keeping the update number fixed. Table 6 presents the final results.
@@ 7.2. Additional Self-refinement during Inference
Table 6. The performance of Generation with Self-refinement during the inference. Specifically, we first report the performance of different number of updated evidences $\alpha$ (Within self-refinement); and then we fix the update number as 1 and iteratively refine the generated response step by step (Multi-step Self-refinement).
@@
7.3. Ablation Study
We investigate the effects of individual steps by providing UniMS-RAG model the ground-truth labels from each step to generate the response, enabling us to analyze and understand the specific effects of each step in a clear and systematic way. Table 7 shows the final results.
7.3.1. Upper bound of UniMS-RAG
First, we note that the inclusion of ground-truth planning labels or retrieval results casts a positive impact on performance. Planning primarily enhances consistency scores, while grounding retrieval results contributes to BLEU1 and Rouge-L scores. The best results are obtained when both signals are combined, serving as the upper bound of our proposed UniMS-RAG.
7.3.2. The effects of evidence attention mask
There are two notable advantages of adding the evidence attention mask: 1) removing the noise of irrelevant evidence to predict the similarity score; and 2) the UniMS-RAG can effectively operate as a retriever, as it directly captures the relationship between individual contexts and corresponding evidence, optimizing its retrieval capabilities. We can clearly find that removing the evidence attention mask leads to performance degradation, especially on KBP dataset since there are more cases that require external sources of knowledge.
7.3.3. Compared with simple planning strategies
We additionally conduct experiments by adopting two simple planning strategies: 1) always use all sources, User-Per and Bot-Per in DuLeMon, and Both in KBP; and 2) always do not use sources, to investigate the effectiveness of the introduced planning step. The results show that the performance is decreased by 63.18%-1.23% on DuLeMon and 42.07%-36.21% on KBP respectively when using all sources. This highlights that indiscriminate use of all sources not only significantly hampers performance when the majority of samples do not necessitate external references (DuLeMon), but also results in a decline even when a substantial portion of samples requires the utilization of sources (KBP). Conversely, a parallel trend is observed when abstaining from using any external sources. When the majority of samples do not require external references, refraining from their usage leads to notable improvement. However, this approach adversely impacts performance when external sources are essential.
Table 7. Ablation study on the impact of different steps and modules in UniMS-RAG.
Model DuLeMon KBP
BLEU1 RougeL P.C BLEU1 RougeL P.C K.C
UniMS-RAG 18.43 20.32 63.18 29.65 32.48 75.51 42.07

### v2

@@ 6 Experimental Results
TABLE IV: The performance (Recall@1) of Retrieval of different types of retrievers. It is note the UniMS-RAG (w/ DPR) means that the similarity labels come from DPR model during training, and for w/ LLMs, the labels comes from zero-shot prompting. 1) DuLeMon: There are 511 samples require to use User-Per and 225 samples require to use Bot-Per; 2) KBP: There are 125 samples require to use PERSONA and 923 samples require to use BOTH.
@@
6.1 Performance of Planning (RQ1)
There are different types of planning decisions in the different datasets: DuLeMon (using NULL, User-Per, Bot-Per and Both sources of knowledge) and KBP (using NULL, Persona, and Both). Table III demonstrates the F1 of planning using different methods under these two datasets. Generally, for prompting-based methods, we can find that two observations: 1) in-context learning (ICL) prompting can not always achieves better performances compared with zero-shot, regardless the dataset and models; and 2) It is not guaranteed that GPT-4 outperforms ChatGPT in all metrics. In detail, GPT4o tends to predict more NULL compared with ChatGPT, especially on DuLeMon dataset. On the other side, for the supervised methods, the SAFARI model performed slightly worse than the BERT model on the DuLeMon dataset, but their performance was comparable on the KBP dataset. Furthermore, the UniMS-RAG achieves best performance on 4 out of 7 metrics. Specifically, for KBP, both UniMS-RAG and SAFARI most frequently predict Both, followed by NULL, and then the Persona case, mirroring the frequency of these cases in the original dataset. However, UniMS-RAG predicts significantly more NULL cases than SAFARI. Some of these NULL predictions by UniMS-RAG are actually Persona in the original dataset, leading to a lower number of Persona predictions by UniMS-RAG. In addition, we also found that the original data distribution has a serious impact on the final planning performance. For example, there are less than 0.1% samples in the training set of DuLeMon requiring Both source of knowledge, resulting in poor performance of all methods at Both. A similar phenomenon is also observed on the KBP dataset. Another reason behind this could be the additional token (i.e., evaluation tokens) introduced during the training stage compared with SAFARI. In general, the planning capability of LLMs still needs to be improved to solve the complex multiple sources planning problem in a dialogue system, particularly when dealing with imbalanced or scarce data sources.
6.2 Performance of Retrieval (RQ2)
To investigate the RQ2, we examine different types of retrievers, including unsupervised methods (BM25, ChatGPT and GPT4o) and supervised methods (RocketQAv2, DPR, and our proposed UniMS-RAG), in order to evaluate the retrieval performance, providing the ground-truth planning labels (except NULL). Table IV presents the Recall@1 (R@1) of the different retrievers.
6.2.1 The performance of baselines
There are several observations can be concluded from the results: 1) GPT4o can achieves better performance than ChatGPT no matter which prompting method is chosen (zero-shot or in-context), and the performance of in-context learning prompting can not always achieves better performance in all datasets. For example, the in-context learning prompting (ICL) performs worse than zero-shot on DuLeMon but better on KBP. We attribute this to the effects of demonstrations and the complexity of different datasets; 2) The overall performance of dense vector (e.g, RocketQAv2 and DPR) mostly is better than ChatGPT and GPT4o, and the performance of ChatGPT is better than sparse retriever (e.g, BM25). The gap of former is bigger than the latter, indicating the large improvement room for current methods using LLMs as retriever, particularly at the context of conversational search; 3) It is observed that DPR performs the best out of these retrievers on KBP datasets while RocketQAv2 performs the best on DuLeMon dataset, revealing the importance of dense retrieval models in this task. We attribute the higher performance of RocketQAv2 on DuLeMon to the similar distribution between DuLeMon dataset and pre-training corpus of RocketQAv2; 4) On KBP datasets, all retrievers performs best at Both-PERSONA and worst at Both-DOCUMENTS, indicating the difficulty to retrieve independent source of knowledge since the semantics between different knowledge from Both-DOCUMENTS are similar to the same underlying persona $p^{*}$, making them more difficult to distinguish. On the other hand, the performance on DuLeMon is even worse than Both-DOCUMENTS on KBP. We suspect this is due to the semantic gap between the dialogue context and used persona in DuLeMon (as shown in Figure 1).
6.2.2 The performance of UniMS-RAG
Besides, we also present the performance of UniMS-RAG as a retriever using different similarity signals (DPR, ChatGPT or GPT4o) during the training. We emphasize here that this kind of functionality of UniMS-RAG as retriever is additional bonus due to the introduction of sub-task: relevance score prediction. It does not mean that we must use itself as retriever during inference in practice. Our main focus is to determine whether UniMS-RAG can be used as a retriever. Additionally, if it can be used as a retriever, we aim to evaluate its performance and potential. The results show that UniMS-RAG can be directly used as a retriever, since the performance of all variants of UniMS-RAG are better than sparse retriever (BM25) and some of them even outperforms some prompting-based methods. In detail, UniMS-RAG w/ DPR even achieves better performance than ChatGPT, revealing the great potential of UniMS-RAG as retriever once we use more high-quality signals (the performance of DPR is better than ChatGPT). Furthermore, UniMS-RAG w/ ChatGPT achieves comparable performance with ChatGPT on DuLeMon and KBP datasets. Since the performance of UniMS-RAG w/ GPT4o is much better than UniMS-RAG w/ ChatGPT due to more accurate similarity labels provided by GPT4o, we believe that UniMS-RAG can distill the capabilities of original retriever as much as possible if we can provide more data and more fine-grained signals. Therefore, we can derive the answer to RQ2 from this analysis: Large Language Models (LLMs) can serve as retrievers directly, achieving comparable performance compared with original retrievers, showcasing a great potential towards a more powerful unified RAG framework.
TABLE V: The performance of Generation of different methods, in which part of the results are directly copied from wang2023large [11]. We follow the official code of SAFARI to conduct evaluation on DuLeMon dataset. We bold the highest performance and underline the second-best performance.
@@ 6.2.2 The performance of UniMS-RAG
6.3 Performance of Generation (RQ3)
@@
To investigate the performance of UniMS-RAG on the response generation task, we conduct two settings: 1) using itself as retriever, which means we use UniMS-RAG to retrieve corresponding evidences from the planned sources of knowledge (i.e, the output of planning step); and 2) using independent retriever (i.e, BM25, DPR, ChatGPT, and GPT4o) to retrieve corresponding evidences from the planned sources of knowledge (i.e, the output of planning step). Table V demonstrates the performance of response generation under both supervised and unsupervised settings.
6.3.1 The performance of baselines
@@ 6.3.2 The performance of UniMS-RAG
Referring to Tabel III and Table IV, the performance of the planning step and retrieval step largely affects the results in the final generation step. Specifically, when using itself as retriever, we can find that UniMS-RAG using signals from DPR (w/ DPR) leads to better performance in contrast to ChatGPT (w/ ChatGPT) or GPT4o (w / GPT4o), revealing the effectiveness of better retriever signals. The results of GPT4o is also slightly better than ChatGPT due to more accurate similarity scores no matter using itself as retriever or using independent retriever. The gap between DPR and ChatGPT on DuLeMon is relatively small since most of cases here do not require the involvement of external sources of knowledge. Thus, we decide to load parameters of UniMS-RAG w/ DPR to conduct evaluation when using independent retriever. In detail, the results again validate the effectiveness of better retriever (w/ DPR $>$ w/ GPT4o $>$ w/ ChatGPT $>$ w/ BM25). We find that the performance gap between different retrievers in DuLeMon dataset is pretty small (less than 1%). We suspect this is highly related to the original data distribution in the test dataset, since most of the samples do not require external persona information. Furthermore, we also find that using independent retriever is mostly better than using itself as retriever except the worst BM25, which is consistent with the findings in the performance of retrieval step. To conclude, it is obvious that our proposed method UniMS-RAG outperforms all other baselines in at least 5 out 7 evaluation metrics no matter using itself as retriever or using independent retriever. Combining the performance of planning, retrieval, and response generation together, we can find that UniMS-RAG is capable of serving as a planner, retriever, and reader in a unified manner, leading to better performance in personalized dialogues.
6.3.3 Discussion of Unified Manner
Despite the clear performance boost from using independent existing retrievers, we emphasize that the performance gain is not solely due to the better relevance scores provided by these retrievers. It is also significantly attributed to our proposed UniMS-RAG, especially the introduction of relevance score prediction task which makes it can attention on relevant evidences while overlooking noisy ones. This conclusion based on two observations: 1) simply using the existing retriever based on SAFARI cannot achieve state-of-the-art results. We have already selected the best-performing version of SAFARI, and all variants of our proposed UniMS-RAG (i.e., using independent retriever and using itself as retriever) largely outperform it, 2) The performance gap between variants of our proposed method is relatively smaller than the gap between ours with baselines, revealing the great potential and flexibility of UniMS-RAG. In practice, we advocate for the unified training of these three sub-tasks in PerDS, as depicted in UniMS-RAG. During inference, the retriever should be selected on a case-by-case basis.
7 Analysis and Discussions
@@
In this section, we choose the best model as shown in our previous experiments to investigate the performance changes under different settings (UniMS-RAG w/ DPR). In detail, we start from the performance of our proposed model with different numbers of retrieved evidence (§7.1), then we study the effects of introduced self-refinement mechanism during the inference stage by re-evaluating the relationship between generated response with the dialogue context or retrieved evidence (§7.2). We present the ablation study to show the effectiveness and rationale of UniMS-RAG (§7.2), followed by the results of human evaluation and case study (See supplementary).
7.1 Different Numbers of Retrieved Results
@@
(a) DuLeMon
(b) KBP
Figure 5: The performance of Generation with different number of retrieved evidences on two datasets.
7.2 Additional Self-refinement during Inference
To investigate the effects of self-refinement during the inferences, we set the number of retrieved evidence as 3 according to the experimental results in the previous section since it leads to the best performance. There are two different ways to refine the generated response: 1) Within self-refinement by providing different number of updated evidences during one-step refinement as shown in Algorithm 1; and 2) Multi-step self-refinement by refining the response step by step while keeping the update number fixed. Table VI presents the final results.
@@ 7.2 Additional Self-refinement during Inference
TABLE VI: The performance of Generation with Self-refinement during the inference. Specifically, we first report the performance of different number of updated evidences $\alpha$ (Within self-refinement); and then we fix the update number as 1 and iteratively refine the generated response step by step (Multi-step Self-refinement).
@@
7.3 Ablation Study
We investigate the effects of individual steps by providing UniMS-RAG model the ground-truth labels from each step to generate the response, enabling us to analyze and understand the specific effects of each step in a clear and systematic way. Table VII shows the final results.
TABLE VII: Ablation study on the impact of different steps and modules in UniMS-RAG.